Grade 11Mechanics

If a ball is thrown vertically upward with speed u, the distance covered during the last time second of its its ascent is

C boy
6 Years agoGrade 11

1 Answer

Vikas TU
6 Years ago
Dear student 
Let body takes T sec to reach maximum height.  
Then v = u - gT  
v = 0, at highest point.  
T=u/g                     ………… (1)
  
Velocity attained by Body in (T - t) sec v = u - g (T -t)  
= u-gT + zt = u-g u/g + gt  
or v = gt                 …………. (2)  
 Distance travelled in last t sec of its ascent  
S = (gt)t -1/2 gt2 = 1/2  gt2